# Mirage of Mastery: Memorization Tricks LLMs into Artificially Inflated Self-Knowledge

### v2

@@ Abstract
When artificial intelligence mistakes memorization for intelligence, it creates a dangerous mirage of reasoning. Existing studies treat memorization and self-knowledge deficits in LLMs as separate issues and do not recognize an intertwining link that degrades the trustworthiness of LLM responses. In our study, we utilize a novel framework to ascertain if LLMs genuinely learn reasoning patterns from training data or merely memorize them to assume competence across problems of similar complexity focused on STEM domains. Our analysis shows a noteworthy problem in generalization: LLMs draw confidence from memorized solutions to infer a higher self-knowledge about their reasoning ability, which manifests as an over 45% inconsistency in feasibility assessments when faced with self-validated, logically coherent task perturbations. This effect is most pronounced in science and medicine domains, which tend to have maximal standardized jargon and problems, further confirming our approach. Significant wavering within the self-knowledge of LLMs also shows flaws in current architectures and training patterns, highlighting the need for techniques that ensure a balanced, consistent stance on models’ perceptions of their own knowledge for maximum AI explainability and trustworthiness. Our code and results are available publicly.  [^1] https://github.com/knowledge-verse-ai/LLM-Memorization_SK_Eval-
Mirage of Mastery: Memorization Tricks LLMs into Artificially Inflated Self-Knowledge
Sahil Kale1thanks: Corresponding author. Email: sahil@k-v.ai  Vijaykant Nadadur1 1Knowledgeverse AI {sahil, vrn}@k-v.ai
@@ 1 Introduction
For true reliability and trustworthiness, AI tools must consistently and accurately recognize the boundary of their capabilities, referred to as self-knowledge (Yin et al., 2023). In the case of large language models (LLMs), it refers to their introspective ability to assess their own cognizance while generating responses (Ni et al., 2024). While methodologies studying LLM self-knowledge are rife (Wang et al., 2023; Wen et al., 2024; Ren et al., 2024), only rarely have other challenges like memorization (Slonski, 2024) and adversarial helpfulness (Ajwani et al., 2024) been attributed as underlying causes for inaccuracies and overconfidence in self-knowledge, a gap needing timely redress.
@@
Recent research highlights two key challenges in large language models: memorisation, the tendency to recall and reproduce training data without true understanding (Satvaty et al., 2025; Prashanth et al., 2024), and a lack of self-knowledge, or the inability to accurately demarcate the limits of their own capabilities (Tan et al., 2024). As data used for LLM training scales to include more problems and solutions, especially in STEM fields, models may conflate the ability of ‘knowing’ solutions with genuine reasoning power to solve such problems. This memorization-driven behaviour raises a critical downstream concern regarding generalization: if LLMs mistake recall for reasoning, they may falsely perceive their self-knowledge to be deeper than it truly is, as shown in Figure 1. Such overconfidence, drawn from memorized problem data and patterns, can lead to unreliable responses, questioning AI trustworthiness.
Figure 2: Methodology to analyze how LLM memorization inflates self-perception using minor task perturbations
Our inspiration can be explained with a simple human parallel: a medical student who memorizes past diagnostic cases may feel highly confident in their ability to diagnose diseases, but without truly understanding the underlying medical principles, they may struggle or misdiagnose slightly altered cases. We propose a novel, universally applicable framework to determine whether LLMs genuinely learn reasoning patterns from training data or merely memorize them, thereby erroneously assuming competence across problems of similar complexity.
In our methodology, we give LLMs the freedom to provide a problem in STEM fields they are confident of solving and evaluate their consistency in feasibility analyses on perturbed problems of similar complexity, requiring the same reasoning patterns. If LLMs rely on memorization for self-knowledge, their accuracy and consistency in answering and determining the feasibility of tasks should falter when faced with minor logical perturbations (Xie et al., 2025). We thus develop a method to identify such inconsistencies in feasibility assessments of logically coherent task variations, revealing if and when LLMs rely on memorization to artificially inflate and generalize their self-knowledge.
@@
Table 1: MIRAGE scores for all models distributed across STEM domains
2 Experimentation Methodology
We present a novel experimentation approach to analyse if LLMs’ tendency to memorise problems and corresponding solutions builds an inflated perception of their own knowledge and capabilities. Our methodology builds on the self-knowledge evaluation technique given by Kale and Nadadur (2025) and the principle from research by Xie et al. (2025) that memorization is characterized by high accuracy on familiar problems but a sharp decline when faced with slight problem variations.
@@ 2 Experimentation Methodology
Table 2: SKEW scores for all models distributed across STEM domains
@@ 2.1 Task Generation
In the initial stage, we use a structured prompt to generate tasks with task instructions and associated data, if any, with the constraint that the LLM finds them completely feasible. During the generation phase, we allow LLMs the flexibility to set their own feasibility boundaries and use the QAP (Yugeswardeenoo et al., 2024) approach while devising prompts to ensure introspection in self-knowledge before providing responses. While we restrict the task domain to STEM fields, no other constraints are imposed in the prompt for task generation (refer to the prompt in Figure 6 in Appendix A). Since LLMs are prone to inconsistencies in self-knowledge assessments (Kale and Nadadur, 2025), we incorporate a separate prompt-based validation step (refer to the prompt in Figure 7 in Appendix A) within the task generation phase to ensure that models are highly confident in their ability to answer the problems they have generated. We feed such validated feasible tasks to the task perturbation pipeline ahead.
@@ 2.2 Task Perturbation Pipeline
Ontology replacement consists of substituting key domain-specific terms in the original task instructions and data with equivalent terminology from the same STEM domain. We also modify relationships between entities in the task while preserving logical coherence in associated data. This step is crucial to minimize terminology-driven memorization while preserving the task’s core formulation and structure. So as to also incorporate basic language level perturbation, we implement this module by setting up a Gemini 1.5 Flash model (Team, 2024) with specific replacement instructions in the prompt shown in Figure 8 in Appendix A.
@@
The data perturbation module uses a simple rule-based approach to introduce an approximately 15 percent variation in all numerical values present in the task data. Since LLMs are prone to memorization of numerical patterns and data (Bordt et al., 2024), this step is crucial to dissociate tasks from such patterns. We also reorder all unordered data elements including lists and arrays in the task data to minimize pattern matching while answering and classifying feasibility of tasks.
@@ 2.3 Task Classification
In the final stage, we feed all perturbed tasks generated through the pipeline to the LLM to attempt (refer to the prompt in Figure 9 in Appendix A). For each task, the LLM is prompted to either generate a conclusive answer (and thus classify the task as feasible) or mark it as infeasible if it finds the task to lie beyond its self-identified reasoning capabilities. Since all tasks are validated to be feasible and are perturbed in a way that maintains domain relevance and complete logical cohesion, an inconsistency in feasibility assessments strongly implies that the LLMs relied on memorization of the data or solution steps of the original task to draw confidence for an inflated sense of self-knowledge and reasoning capacity in that domain.
@@ 3.2 Setup
For comprehensive analysis, we experiment with a wide range of high-performance models. Since our methodology is applicable across both closed and open-source models, we choose 2 closed-source models, GPT-4o (OpenAI, 2024) and Claude 3.7 Sonnet (Anthropic, 2024), and 2 open-source models, Mistral Large 24.11 (AI, 2024) and DeepSeek v3 (DeepSeek-AI, 2024), in our evaluation. For all models, detailed model parameters are provided in Section A.2 in Appendix A. For each STEM domain and model, we generate 34 original tasks and perturb each task 3 times, leading to a dataset of 102 perturbed tasks per domain and 408 across all domains. MIRAGE scores across LLMs for all domains are presented in Table 1, while SKEW values analysing self-knowledge inconsistencies are given in Table 2.
@@ 4 Quantifying Memorization-inflated Self-Knowledge
Self-knowledge inflation due to memorization is striking. The consistently high MIRAGE scores for all models, as shown in Table 1 and visualised in Figure 3, point to an alarming fundamental flaw: models are not just mistaking recall for reasoning but also inflating their self-knowledge perception using this memorization-driven confidence. Even the best high-performance models like GPT-4o and Mistral Large 24.11 change their feasibility stance about slightly perturbed tasks over 45% of time, meaning that LLMs are systematically overestimating their ability based on memorized solutions and data rather than true reasoning awareness. A likely cause of this widespread memorization-induced overconfidence could be the prevalence of STEM benchmarks and evaluations in training data, which may act more towards reinforcing memorization patterns rather than fostering robust reasoning in LLMs. This finding highlights a significant trust issue and a critical vulnerability in AI reliability and generalization.
@@ 5 Impact of Self-Knowledge Inconsistencies
LLMs are not trustworthy enough for critical real-world applications. Our findings highlight fundamental challenges and potential advancements in deploying LLMs for high-stakes domains like healthcare, law, and research, where memorized or unreliable reasoning can cause critical real-world consequences. We demonstrate how different LLMs exhibit domain-specific weaknesses across distinct STEM fields and recommend adaptive LLM routing strategies to be wary of these issues during the selection of models for important tasks. Given that most models exhibit artificially inflated self-knowledge in science and medicine due to memorization, we emphasize the need for cautious deployment in these fields. Implementing safeguards like confidence thresholds and source markers can help flag uncertain responses, ensuring users are aware of potential inaccuracies before using AI-generated outputs elsewhere. Due to their inflated self-perception, models are prone to generating responses even when lacking sufficient knowledge, rather than abstaining. Hence, human-in-the-loop fallback strategies are still important in LLM-powered applications for maximum trustworthiness. We urge future research to prioritize methods that enhance self-knowledge robustness, reduce overreliance on memorization, and establish mechanisms for models to recognize and convey uncertainty in their own knowledge.
@@ 6 Related Work
Memorization in LLMs: Memorization, be it verbatim or contextual, has garnered significant attention due to its implications for privacy, trustworthiness, and the generalization capabilities of models. Hartmann et al. (2023) proposed a taxonomy for memorization in LLMs, which was further improved along dimensions of granularity, retrievability, and desirability by Satvaty et al. (2025). Based on such classifications, numerous techniques to uncover such behaviour have been presented including methods that utilize injected sequences (Huang et al., 2024), neuron activations (Slonski, 2024), counterfactual reasoning tasks (McCoy et al., 2024) or dynamic, prefix-dependent soft prompts (Wang et al., 2024). Memorization analyses have also included multimodal behaviour (Ju et al., 2025) and fine-tuned models (Zeng et al., 2024) in recent studies.
Self-knowledge in LLMs: Most studies on self-knowledge in LLMs assess responses in fixed question-answering tasks, equating the self-knowledge problem to binary classifications of answerability (Zhang et al., 2024; Ren et al., 2024; Wen et al., 2024). Recent studies have diversified to understand self-cognition (Chen et al., 2024), self-recognition using security questions (Davidson et al., 2024), embedding interpretation (Speicher et al., 2024), and generation-validation consistency of LLMs (Li et al., 2023) from a self-knowledge perspective. However, most uncertainty detection methods currently lack feasible alternatives for closed-source models where intermediate stated cannot be analysed (Yona et al., 2024). Our research builds on previous findings to establish a link between self-knowledge overconfidence and another persistent challenge of memorization in LLMs, a connection that, to our knowledge, has not been explored before.
@@ 7 Conclusion
Overconfidence in capabilities is a severe problem in LLMs, hampering their trustworthiness. In this research, we show how models very likely draw confidence from ‘knowing’ solutions to develop an inflated perception of their reasoning power. We use a novel experimentation approach towards this goal by generating feasible tasks and checking if slight problem perturbations change self-knowledge assessments to analyse if LLM’s tendency to memorise data actually causes overreach in self-knowledge.
@@
    Small sample size: Since our experiments rely on LLM-generated tasks, which may eventually repeat, we use just over 400 tasks per model to ensure diversity and distinctiveness, although the size may be perceived as relatively small compared to other studies. We plan to conduct more exhaustive testing on more models, too, in future work.
@@
Appendix A Appendix
A.1 Prompt formats
@@
A.2 Model parameters
We report the generation parameters for all models used in our experiments to ensure transparency and reproducibility. All models were accessed through provider APIs, and common parameter values set during task generation and task classification are listed in Tables 3 and 4, respectively. The model sizes of all closed-source models are approximate and taken from Abacha et al. (2025).
OpenAI Models: We run our experiments on the flagship models, GPT-4o (~200B parameters). We use the OpenAI Python SDK to access the models via API, specifying seed=1234 and n=1 along with the parameter values listed in the respective tables. All other parameters are kept to default values.
@@ A.2 Model parameters
Table 3: Parameters used across all models during task generation

### v3

@@ Abstract
When artificial intelligence mistakes memorization for intelligence, it creates a dangerous mirage of reasoning. Existing studies treat memorization and self-knowledge deficits in large language models (LLMs) as separate issues and do not recognize an intertwining link that degrades the safety and trustworthiness of LLM responses. In our study, we utilize a novel framework to ascertain if LLMs genuinely learn reasoning patterns from training data or merely memorize them to assume competence across problems of similar complexity focused on STEM domains. Our analysis shows a noteworthy problem in generalization: LLMs draw confidence from memorized solutions to infer a higher self-knowledge about their reasoning ability, which manifests as an over 45% inconsistency in feasibility assessments when faced with validated, logically coherent task perturbations. This effect is most pronounced in science and medicine domains, which tend to have maximal standardized jargon and problems, further confirming our approach. Such overconfidence poses critical AI safety risks in trust-sensitive domains, like healthcare, law, and scientific research, where unreliable responses can have severe consequences. Significant wavering within the self-knowledge of LLMs also shows flaws in current architectures and training patterns, highlighting the need for techniques that ensure a balanced, consistent stance on models’ perceptions of their own knowledge. Our code and results are available publicly.  [^1] https://github.com/Sahil-R-Kale/mirage_of_mastery
@@ 1 Introduction
For true reliability and trustworthiness, AI tools must consistently and accurately recognize the boundary of their own capabilities, referred to as self-knowledge [yin2023largelanguagemodelsknow]. In the case of large language models (LLMs), it refers to their introspective ability to determine which tasks and prompts are feasible precisely and with certainty. [ni2024llmsneedretrievalaugmentation, kale2025linedutyevaluatingllm]. Previous methodologies studying LLM self-knowledge are rife [wang-etal-2023-self-knowledge, wen2024perceptionknowledgeboundarylarge, ren2024investigatingfactualknowledgeboundary], and have shown that LLMs lack a grounded understanding of their own knowledge boundaries. However, only rarely have other challenges like memorization [slonski2024detectingmemorizationlargelanguage] and adversarial helpfulness [ajwani2024llmgeneratedblackboxexplanationsadversarially] been attributed as underlying causes for inaccuracies and overconfidence in self-knowledge, a gap needing timely redress.
@@
Recent research highlights two key challenges in large language models: memorisation, the tendency to recall and reproduce training data without true understanding [satvaty2025undesirablememorizationlargelanguage, prashanth2024recitereconstructrecollectmemorization], and, as previously described, a lack of self-knowledge, or the inability to accurately demarcate the limits of their own capabilities [tan2024iunderstandicreate]. As data used for LLM training scales to include more problems and solutions, especially in STEM fields, we believe models may conflate the ability of ‘knowing’ solutions with genuine reasoning power to solve such problems. This memorization-driven behaviour raises a critical downstream concern regarding generalization: if LLMs mistake recall for reasoning, they may falsely perceive their self-knowledge to be deeper than it truly is, as shown in Figure 1. Such overconfidence, drawn from memorized problem data and patterns, can lead to unreliable responses, harming AI trustworthiness and safety.
Our inspiration can be explained with a simple human parallel: a medical student who successfully diagnoses and treats one patient will not assume they can automatically solve any other case of similar difficulty without understanding the underlying medical principles. In contrast, we find that LLMs tend to behave as if they do, drawing undue confidence from memorized examples. We propose a novel, universally applicable framework to determine whether LLMs genuinely learn reasoning patterns from training data or merely memorize them, thereby erroneously assuming competence across problems of similar complexity.
In our methodology, we give LLMs the freedom to provide a problem in STEM fields they are confident of solving and evaluate their consistency in feasibility analyses on perturbed problems of similar complexity, requiring the same reasoning patterns. If LLMs rely on memorization for self-knowledge, their accuracy and consistency in answering and determining the feasibility of tasks should falter when faced with minor logical perturbations [xie2025memorizationlargelanguagemodels]. We thus develop a method to identify such inconsistencies in feasibility assessments of logically coherent task variations, revealing if and when LLMs rely on memorization to artificially inflate and generalize their self-knowledge.
@@
Figure 2: Methodology to analyse how LLM memorization inflates self-perception using minor task perturbations
2 Experimentation Methodology
We present a novel experimentation approach to analyse if LLMs’ tendency to memorise problems and corresponding solutions builds an inflated perception of their own knowledge and capabilities. Our methodology builds on the self-knowledge evaluation technique given by [kale2025linedutyevaluatingllm] and the principle from research by [xie2025memorizationlargelanguagemodels] that memorization is characterized by high accuracy on familiar problems but a sharp decline when faced with slight problem variations.
Table 1: MIRAGE scores for all models distributed across STEM domains
@@ 2 Experimentation Methodology
Table 2: SKEW scores for all models distributed across STEM domains
@@ 2.1 Task Generation
In the initial stage, we use a structured prompt to generate tasks with task instructions and associated data, if any, with the constraint that the LLM finds them completely feasible. During the generation phase, we allow LLMs the flexibility to set their own feasibility boundaries and use the QAP [yugeswardeenoo-etal-2024-question] approach while devising prompts to ensure introspection in self-knowledge before providing responses. While we restrict the task domain to STEM fields, no other constraints are imposed in the prompt for task generation (refer to the prompt in Figure 6 in Appendix Appendix). Since LLMs are prone to inconsistencies in self-knowledge assessments [kale2025linedutyevaluatingllm], we incorporate a separate prompt-based validation step (refer to the prompt in Figure 7 in Appendix Appendix) within the task generation phase to ensure that models are highly confident in their ability to answer the problems they have generated. Also, we conduct spot checks using STEM field experts to ensure task feasibility. We feed such validated feasible tasks to the task perturbation pipeline ahead.
@@ 2.2 Task Perturbation Pipeline
Ontology replacement consists of substituting key domain-specific terms in the original task instructions and data with equivalent terminology from the same STEM domain. We also modify relationships between entities in the task while preserving logical coherence in associated data. This step is crucial to minimize terminology-driven memorization while preserving the task’s core formulation and structure. So as to also incorporate basic language level perturbation, we implement this module by setting up a Gemini 1.5 Flash model [geminiteam2024gemini15unlockingmultimodal] with specific replacement instructions in the prompt shown in Figure 8 in Appendix Appendix.
@@
The data perturbation module uses a simple rule-based approach to introduce an approximately 15 percent variation in all numerical values present in the task data. Since LLMs are prone to memorization of numerical patterns and data [bordt2024elephantsforgetmemorizationlearning], this step is crucial to dissociate tasks from such patterns. We also reorder all unordered data elements including lists and arrays in the task data to minimize pattern matching while answering and classifying feasibility of tasks.
For all perturbed tasks, to ensure feasibility, we take the help of STEM field experts to manually verify feasibility and remove samples where difficulty level is increased due to language vocabulary limitations, or data-related inconsistencies.
@@ 2.3 Task Classification
In the final stage, we feed all perturbed tasks generated through the pipeline to the LLM to attempt (refer to the prompt in Figure 9 in Appendix Appendix). For each task, the LLM is prompted to either generate a conclusive answer (and thus classify the task as feasible) or mark it as infeasible if it finds the task to lie beyond its self-identified reasoning capabilities. Since all tasks are validated to be feasible and are perturbed in a way that maintains domain relevance and complete logical cohesion, an inconsistency in feasibility assessments strongly implies that the LLMs relied on memorization of the data or solution steps of the original task to draw confidence for an inflated sense of self-knowledge and reasoning capacity in that domain.
@@ 3.1 Formulation and Metrics
Owing to the novelty of exploring such a relation between known LLM problems, and a lack of suitable existing benchmarks, we devise two new metrics based on simple averages to quantify our experimental results.
@@ 3.2 Setup
For comprehensive analysis, we experiment with a wide range of high-performance models. Since our methodology is applicable across both closed and open-source models, we choose 2 closed-source models, GPT-4o [openai2025gpt4o] and Claude 3.7 Sonnet [anthropic2025claude3], and 2 open-source models, Mistral Large 24.11 [mistral2025large2407] and DeepSeek v3 [deepseekai2024deepseekv3], in our evaluation. For all models, detailed model parameters are provided in Section .2 in Appendix Appendix. For each STEM domain and model, we generate 34 original tasks and perturb each task 3 times, leading to a dataset of 102 perturbed tasks per domain and 408 across all domains. MIRAGE scores across LLMs for all domains are presented in Table 1, while SKEW values analysing self-knowledge inconsistencies are given in Table 2.
@@ 4 Quantifying Memorization-inflated Self-Knowledge
Self-knowledge inflation due to memorization is striking. The consistently high MIRAGE scores for all models, as shown in Table 1 and visualized in Figure 3, point to an alarming fundamental flaw: models are not just mistaking recall for reasoning but also inflating their self-knowledge perception using this memorization-driven confidence. Even the best high-performance models like GPT-4o and Mistral Large 24.11 change their feasibility stance about slightly perturbed tasks over 45% of time, meaning that LLMs are systematically overestimating their ability based on memorized solutions and data rather than true reasoning awareness. A likely cause of this widespread memorization-induced overconfidence could be the prevalence of STEM benchmarks and evaluations in training data, which may act more towards reinforcing memorization patterns rather than fostering robust reasoning in LLMs. This finding highlights a significant trust issue and a critical vulnerability in AI reliability and generalization.
@@ 5 Impact of Self-Knowledge Inconsistencies
LLMs are not trustworthy enough for critical real-world applications. Our findings highlight fundamental challenges and potential advancements in deploying LLMs for high-stakes domains like healthcare, law, and research, where memorized or unreliable reasoning can cause critical real-world consequences. We demonstrate how different LLMs exhibit domain-specific weaknesses across distinct STEM fields and recommend adaptive LLM routing strategies [ong2024routellmlearningroutellms] to be wary of these issues during the selection of models for important tasks. Given that most models exhibit artificially inflated self-knowledge in science and medicine due to memorization, we emphasize the need for cautious deployment in these fields. Implementing safeguards like confidence thresholds and source markers can help flag uncertain responses, ensuring users are aware of potential inaccuracies before using AI-generated outputs elsewhere. Due to their inflated self-perception, models are prone to generating responses even when lacking sufficient knowledge, rather than abstaining. Hence, human-in-the-loop fallback strategies are still important in LLM-powered applications for maximum trustworthiness. We urge future research to prioritize methods that enhance self-knowledge robustness, reduce over-reliance on memorization, and establish mechanisms for models to recognize and convey lack in their own knowledge.
@@ 6 Related Work
Memorization in LLMs: Memorization, be it verbatim or contextual, has garnered significant attention due to its implications for privacy, trustworthiness, and the generalization capabilities of models. [hartmann2023sokmemorizationgeneralpurposelarge] proposed a taxonomy for memorization in LLMs, which was further improved along dimensions of granularity, retrievability, and desirability by [satvaty2025undesirablememorizationlargelanguage]. Based on such classifications, numerous techniques to uncover such behaviour have been presented including methods that utilize injected sequences [huang-etal-2024-demystifying], neuron activations [slonski2024detectingmemorizationlargelanguage], counterfactual reasoning tasks [McCoy2024Embers] or dynamic, prefix-dependent soft prompts [wang-etal-2024-unlocking]. Memorization analyses have also included multimodal behaviour [ju2025watchalbuminadvertentprivacy] and fine-tuned models [zeng-etal-2024-exploring] in recent studies.
Self-knowledge in LLMs: Most studies on self-knowledge in LLMs assess responses in fixed question-answering tasks, equating the self-knowledge problem to binary classifications of answerability [zhang2024largelanguagemodelsgood, ren2024investigatingfactualknowledgeboundary, wen2024perceptionknowledgeboundarylarge]. Recent studies have diversified to understand self-cognition [chen2024selfcognitionlargelanguagemodels], self-recognition using security questions [davidson2024selfrecognitionlanguagemodels], embedding interpretation [speicher2024understandingmemorisationllmsdynamics], and generation-validation consistency of LLMs [li2023benchmarkingimprovinggeneratorvalidatorconsistency] from a self-knowledge perspective. However, most existing methods currently lack feasible alternatives for closed-source models where intermediate stated cannot be analysed [yona-etal-2024-large]. Our research builds on previous findings to establish a link between self-knowledge overconfidence and another persistent challenge of memorization in LLMs, a connection that, to our knowledge, has not been explored before.
@@ 7 Conclusion
Overconfidence in capabilities is a severe problem in LLMs, hampering their trustworthiness. In this research, we show how models very likely draw confidence from ‘knowing’ solutions to develop an inflated perception of their reasoning power. We use a novel experimentation approach towards this goal by generating feasible tasks and checking if slight problem perturbations change self-knowledge assessments to analyse if LLM’s tendency to memorize data actually causes overreach in self-knowledge.
@@
    Small sample size and custom metrics: Since our experiments rely on LLM-generated tasks, which may eventually repeat, we use just over 400 tasks per model to ensure diversity and distinctiveness, although the size may be perceived as relatively small compared to other studies. Similarly, our metrics are defined anew owing to the uniqueness of the exploration. We plan to conduct more exhaustive testing on more models, too, in future work. Also, we plan to establish standard benchmarks and metrics for such evaluation.
@@
Appendix
.1 Prompt formats
@@
.2 Model parameters
We report the generation parameters for all models used in our experiments to ensure transparency and reproducibility. All models were accessed through provider APIs, and common parameter values set during task generation and task classification are listed in Tables 3 and 4, respectively. The model sizes of all closed-source models are approximate and taken from [abacha2025medecbenchmarkmedicalerror].
OpenAI Models: We run our experiments on the flagship models, GPT-4o (~200B parameters). We use the OpenAI Python SDK to access the models via API, specifying seed=1234| and
@@
Table 3: Parameters used across all models during task generation